In March, 2017, Mayton Mining Co. purchased a coal mine for $12,000,000. Total possible coal to be mined is estimated at 2,000,000 tons. Mayton is required by law to restore the land to a reasonable condition after the conclusion of mining operations at an estimated cost of $750,000. Mayton estimates the land will then be worth $2,000,000. The company incurred $2,800,000 of development costs preparing the mine for production. During 2017, 400,000 tons were removed and 310,000 tons were sold.

Calculate the depletion that Mayton had in 2017 and prepare the depletion journal entry. Also prepare the entry to COGS for the sales during 2017.
Answer #1

Coal mine asset cost that will be depleted = 12000000 + 2800000 + 750000 - 2000000 = 13550000

Depletion per ton of extraction = 13550000 / 2000000 = $ 6.775 per ton

Depletion for 400000 tons removal = 6.775 * 400000 = $ 2710000

Depletion entry:

Coal account Debit 2710000

Coal mine asset account Credit 2710000

Cost of goods sold entry:

Cost of goods sold Debit 2100250

Coal account Credit 2100250